OpenStudy (anonymous):

please help me jason paid 16.88 for a 6-lb mixture of granola and dried apple chucks. If the granola cost 2.10 per pound and the dried apple chunks cost 4.24 per pound, how much pounds of each did he buy?

OpenStudy (accessdenied):

Write two equations that model the information given and then find the solution of them. Consider these as variables: \( g = \text{lbs of granola} \\ d = \text{lbs of dried apple chunks}\) You need an equation for the total # of lbs and the total cost.

OpenStudy (anonymous):

so how would I write it out to solve it

OpenStudy (accessdenied):

Well, we have the unknowns specified. We know that the sum of those should amount to the \(6\ \text{lb} \) mixture we have. We also know that we spent \($16.88\) total, where the granola costs \($2.10\ \text{per lb}\) and the apple chunks cost \($4.24\ \text{per lb}\). The amount we spend for \(g\ \text{lbs}\) of granola should then be \(2.10g\) and the same done with \(d\ \text{lbs}\) of dried apple chunks.

OpenStudy (anonymous):

8.44 pounds of each

OpenStudy (accessdenied):

8.44 would be much larger than the sum we were given, 6 lbs total.

OpenStudy (accessdenied):

Modelling the data given, I find the equations: \(g + d = 6\), and \(2.10g + 4.24d = 16.88\). because "\(g + d\)" represents the total amount we bought (\(6\ \text{lbs}\)), and "\(2.10g + 4.24d\)" gives us the total / sum of what we spent, which should be \($16.88\). Makes sense?

OpenStudy (anonymous):

iam trying to figure out the answer now cause iam still a little confuse

OpenStudy (anonymous):

i should let x be the part i dnt know right

OpenStudy (accessdenied):

Uhh, we can use any variable to represent what we don't know. I used \(g\) and \(d\) because they just help remember what we're looking for while working. You could also use \(x\) and \(y\) if you wanted to, though.

OpenStudy (accessdenied):

Like, if we used \(x\) for \(\text{lbs of granola}\) and \(y\) for \(\text{lbs of dried apple chunks}\), the equations would look like this: \(x + y = 6\) \(2.10x + 4.24y = 16.88\)

OpenStudy (accessdenied):

The variable is merely a representation of what we don't know and need to find. The name we give it does not affect the problem, it's just there so we can actually play with the equation more easily than with stuff like boxes. :)

OpenStudy (anonymous):

alright so the equation is 2.10+ 4.24 =16.88 but what about the 6lbs

OpenStudy (anonymous):

i understand what u r saying i jus dnt get how i suppose to set it up to solve

OpenStudy (accessdenied):

x + y = 6 < total lbs eq. 2.10x + 4.24y = 16.88 < total cost eq. I think the easiest way to solve this system of equations is just substitution. x + y = 6 x = 6 - y 2.10(6 - y) + 4.24y = 16.88 Like that (though it might help to multiply both sides of that by 100 to make the actual arithmetic easier since decimals are usually yucky)

OpenStudy (anonymous):

that seem so confuseing but i will try to solve it jus like that

OpenStudy (accessdenied):

Hmm.. could you explain whats confusing about it? We're basically just using properties of equality (Adding stuff to both sides) so we can get a one-variable equation that allows us to solve for one variable, and then getting the other variable using the first variable we found.

OpenStudy (anonymous):

my answer dnt seem right I take 4.28 divide it into 6.34 which is .68

OpenStudy (accessdenied):

Could you try writing all the work? I'll attempt to identify what went wrong.

OpenStudy (anonymous):

2.10(6 - y) + 4.24y = 16.88 I took 2.10 and times it by 6 which is 12.6 then 2.10 times y which is 2.10y so the new equation is 12.6-2.10y+4.24y=16.88 is that right so far

OpenStudy (accessdenied):

Yup, that's correct.

OpenStudy (anonymous):

then i put the variables togather 10y +4.24y=14.24 so now i got 12.6-14.24y=16.88 then i have to get the y by it self so i substracted 12.6 to both side which leaves 14.24y=4.28 then divid 14.24 to both side leaves y=3.33

OpenStudy (anonymous):

or is it jus y=3.

OpenStudy (accessdenied):

When you're putting the variables together, you should add together the terms with the "y". -2.10y + 4.24y = 2.14y So: 12.6 + 2.14y = 16.88

OpenStudy (anonymous):

thank u

OpenStudy (accessdenied):

So, you should get y=2 with that Then we have to go back to the equation with x and put in that y-value: x = 6-y y=2 x = 6-2 x = 4

OpenStudy (accessdenied):

The moral of the story, I suppose, is that: Algebra is all about using relations of these symbols that represent values. Once we have some relation (like, x+y=6), we can use those basic properties of Math like addition/multiplication/etc. to both sides For example, if we know that x+y=6, then we can subtract y from both sides because of addition property of equality. x=6-y. When we know that second equation (2.10x + 4.24y = 16.88), we can also use that information about x=6-y since they're the \(same\) variables, just in a different relation. I hope that makes sense and I have helped you. :)

OpenStudy (accessdenied):

Sorry, I should say, "subtraction property of equality", but it works the same.

OpenStudy (anonymous):

i had to read it a few times but i think i got it thanks

OpenStudy (accessdenied):

You're welcome. It's a really basic idea that just goes so far in Mathematics in general (and usually stumps people easily, in my classes). :P Once you know the equation, its basically all just using properties to derive things. Creating the equations just follows from stating the things you need to know, then figuring out how they are related in a problem. :)

OpenStudy (anonymous):

lol

OpenStudy (accessdenied):

Not sure what the 'lol' is for.